There is a long straight wire carrying current I = 497 amperes. At a distance of 1.4 meter away, a small piece of wire (length 3 cm) carries a current of 15 amperes in a direction parallel to the long wire.

(a) Calculate the value of the B field generated by the long wire at the position of the small wire.

B = ?T     

(b) Using your result for B just found, calculate the force on the small wire.

F = ?N     
Answer #1

B = u0 * i / (2*pi*0.7) = 4*pi*10-7 * 497 / (2*pi*1.4) = 7.10*10-5 T

F = B*i*L = (7.10*10-5 T )*15*0.03 = 3.19*10-5 N